OpenStudy (anonymous):

Compute.\[\lim_{n\to\infty} \left(\frac{n^n}{n!}\right)^{1/n}\]

OpenStudy (anonymous):

\[\lim_{n\to \infty}\frac n{(n!)^{1/n}}\]Can I do this?

OpenStudy (experimentx):

inf^0 on the bottom

OpenStudy (anonymous):

\[\large \lim_{n\to \infty}e^{\ln n - \frac1n\ln n!}\]

OpenStudy (experimentx):

i was thinking the same.

OpenStudy (experimentx):

I've got a bad feeling about this .... does it converge??

OpenStudy (anonymous):

But I don't know how to get it from here :/ We can't even use l'hospital.

OpenStudy (experimentx):

ln n! = ln n + ln n-1 + ln n-2 + ....

OpenStudy (anonymous):

The answer is e, I think.

OpenStudy (experimentx):

Yes ... it converges

OpenStudy (experimentx):

ans 0

OpenStudy (experimentx):

No ... i think i forgot to multiply by n.

OpenStudy (experimentx):

let's ask wolf bro first.

OpenStudy (experimentx):

wolf bro says e

OpenStudy (anonymous):

\[\frac n{(n!)^{1/n}} = n\cdot\frac{1}{n^{1/n}}\cdot \frac{1}{(n-1)^{1/n}}\cdots \frac{1}{2^{1/n}}\cdot\frac{1}{1^{1/n}}\]

OpenStudy (anonymous):

Hmm or maybe \[\left(\frac{n}{n-1}\cdot \frac n {n-2}\cdots\frac n 2 \cdot\frac n1\right)^{1/n}\]

OpenStudy (experimentx):

yes ... that might be

OpenStudy (experimentx):

let's put it in ln on the top of e

OpenStudy (anonymous):

\[\huge e^{\frac1n \ln\left(\frac{n}{n-1}\cdot \frac n {n-2}\cdots\frac n 2 \cdot\frac n1\right)}\]

OpenStudy (anonymous):

I wish I knew more math other than the fancy tex.

OpenStudy (experimentx):

expand ln and ln(n/n) + ln(n/n-1) ...

OpenStudy (experimentx):

and not n-1 it's n at the beginning.

OpenStudy (anonymous):

But we can cancel, can't we? ln1 = 0

OpenStudy (experimentx):

yes .. only first one.

OpenStudy (anonymous):

1/n(ln n/(n-1) + ln n/(n-2) + ... + ln (n/1))

OpenStudy (experimentx):

1/n(ln n/(n) + ln n/(n-1) + ... )

OpenStudy (anonymous):

But ln n -> infity, maybe we can do something with that 1/n

OpenStudy (experimentx):

(ln n - ln(n-1))/n + (ln n - ln(n-2))/n + ...

OpenStudy (anonymous):

It's the same as before

OpenStudy (experimentx):

Let's use L'hospital rule here 1/n + (ln n - ln(n-1))/n + (ln n - ln(n-2))/n + ...

OpenStudy (experimentx):

(ln n - ln(n-1))/n = (1/n - 1/(n-1)) / 1 + (1/n - 1/(n-2)) + ...

OpenStudy (experimentx):

stuck ... where does out one come from??

OpenStudy (anonymous):

Oh hmm Nice 1/n - 1/(n-1) + 1/n - 1/(n-2) + ... + 1/n - 1 1 - (1/(n-1) + 1/(n-2) + ... + 1/(2) + 1)

OpenStudy (anonymous):

-(1/(n-1) + 1/(n-2) + ... 1/2)

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

hmm i feel i did something wrong

OpenStudy (anonymous):

maybe @eliassaab can help.

OpenStudy (experimentx):

i am out of clue.

OpenStudy (anonymous):

I will try searching on mathstackexchange someone must have asked this question.

OpenStudy (experimentx):

that would spoil the fun.

OpenStudy (experimentx):

LOL

OpenStudy (experimentx):

I am getting answer zero.

OpenStudy (anonymous):

yeah :/ hmm then i will bump the question.

OpenStudy (experimentx):

e^0

OpenStudy (experimentx):

wolf bro says this must be http://www.wolframalpha.com/input/?i=lim+n-%3Einf+1%2Fn+ln%28n%5En%2Fn%21%29

OpenStudy (anonymous):

\[\left(\frac{n}{n-1}\cdot \frac n {n-2}\cdots\frac n 2 \cdot\frac n1\right)^{1/n}= \left(\frac1{1 - 1/n} \cdot \frac1 {1-2/n}\cdots\right)^{1/n}\]

OpenStudy (experimentx):

ln 1 = 0 ,,,, damn.

OpenStudy (experimentx):

i think i understand the error of my method.

OpenStudy (experimentx):

I am out of clues.

OpenStudy (experimentx):

the left would be infinity.

OpenStudy (anonymous):

\[\Large e^{\ln n - \frac{\ln n!}n} =\frac {e^{\ln n} }{e^{-\frac{\ln n!}n}}\]

OpenStudy (anonymous):

Sorry. hehe

OpenStudy (anonymous):

What am I doing!?!?!

OpenStudy (experimentx):

the top would be infinity and the bottom would be zero.

OpenStudy (experimentx):

Ideas generate like that.

OpenStudy (anonymous):

\[\Large e^{\ln n - \frac{\ln n!}n} =\frac {e^{\ln n} }{e^{\frac{\ln n!}n}}\]Hmm are you sure bottom would be infinity n! >>> n

OpenStudy (anonymous):

Can't we apply l'hospital on the bottom part now? maybe somehow we will get e^(ln n-1) hehe

OpenStudy (anonymous):

I don't think we can :/

OpenStudy (experimentx):

the bottom part woulbe be e^1

OpenStudy (anonymous):

The problem is with n! at first it seems like a perfect variable but at the end it gets constant, and then you wonder what happens in between.

OpenStudy (experimentx):

yeah ... that's why my method is not working. and n -something = n/some no

OpenStudy (anonymous):

@phi @nikvist

OpenStudy (anonymous):

@mr.math

OpenStudy (anonymous):

(a) Show, using upper and lower Rieman Sums that \[ \int_1^n \ln(x) \le \sum_{k=2}^n \ln (k) \le \int_1^n \ln(x) + \ln(n) \] (b) Compute \[ \int_1^n \ln(x) \] Use (a) and (b) to prove \[ (c) \quad e \left ( \frac n e \right)^n\le n! \le e \left ( \frac n e \right)^n n \] use (c) to show that the sequence in question tends to e.

OpenStudy (experimentx):

I think you can get the answer from third post ...

OpenStudy (anonymous):

Hmm third post?

OpenStudy (experimentx):

limn→∞ e^(lnn−1/n *lnn!) substituting what you done before will give the answer. Well I still like approximation way. search for Srinivasa Ramanujan http://en.wikipedia.org/wiki/Factorial

OpenStudy (anonymous):

ohh hmm i will search for him, thanks.